# Can Large Language Models Understand Symbolic Graphics Programs?

### v3

@@ 1 Introduction
  • To improve the symbolic program understanding, we collect an instruction-following dataset and propose a new finetuning method, called symbolic instruction tuning.
  • We introduce a symbolic MNIST dataset, where the symbolic program understanding can be challenging so that even a powerful LLM (i.e., GPT-4o) can only obtain nearly a chance-level performance while the rendered image remains very easy to recognize.
  • Finally, we find that SIT can also improve general reasoning performance.
@@ 3 Why is Understanding Symbolic
To showcase why semantic understanding of symbolic graphics problems require multiple sophisticated reasoning abilities, we provide a few qualitative examples of LLM’s output in Figure 3 and Figure 4. Figure 3 shows a qualitative example of how OpenAI-o1 reasons over the CAD program. The reasoning process is highly nontrivial, as it requires multiple reasoning abilities, such as numeric perception, spatial reasoning, geometric understanding, long-range planning and common sense. This is also part of the reason that SGP-Bench can well evaluate the general reasoning ability of LLMs. In Figure 4, we query different LLMs (from weak to strong: Llama-3.1-8B, Llama-3.1-70B, OpenAI-o1) by asking which digit this given SVG program represents. We can observe that all LLMs start to reason from the low-level curves and gradually build up its understanding from local components to semantic elements. Specifically, we can observe that LLMs understand the symbolic graphics program through line-by-line reasoning and then combine the detailed reasoning into an overall semantic understanding. This process is particularly intriguing because it shows that LLMs understand symbolic programs through actual reasoning rather than memorization. More interestingly, the more powerful LLMs, e.g., OpenAI-o1, show better general understanding of the symbolic graphics program and fine-grained grounding of the program, which is consistent with the results on SGP-Bench.
@@
To highlight the importance of symbolic graphics programs, we give another example in Figure 5, where a powerful LLM like OpenAI-o1 can suffer from the visual spurious correlation, and in contrast, this spurious correlation can be avoided when using the symbolic graphics program as the visual representation. Specifically, we construct a visual example that resembles the Ebbinghaus illusion, but the conclusion differs from the classic Ebbinghaus illusion (i.e., two orange circles are of the same size, although they look different) as we intentionally make one of the orange circle obviously larger than the other (shown in Figure 5). Then we feed this curated example to OpenAI-o1 and ask which orange circle is larger. We compare three cases: (a) image input; (b) symbolic program input; and (c) image input with indirect symbolic program. As long as we first ask the LLM to translate the image into symbolic program, the LLM will no longer suffer from the spurious correlation. The same phenomenon also appears for the Müller-Lyer illusion using Claude-3.5-Sonnet (Figure 6). These examples validate the advantages of reasoning the visual world through symbolic graphics programs.
@@ 4.2 Benchmarking Semantic Understanding
SVG dataset statistics. We collect in total $1,085$ SVG programs covering $19$ categories, and each program has $4$ semantic multiple-choice questions (with $4$ options), resulting in a total of $4,340$ questions. We make sure that answers are evenly distributed across $4$ options. Detailed dataset statistics are in Figure 8(a). Our SVG benchmark consists of $5$ different types of questions, including “Semantic”: $1,085$ questions, “Color”: $864$ questions, “Shape”: $1,217$ questions, “Count”: $819$ questions, and “Reasoning”: $355$ questions. “Semantic” tests the global semantic meaning of the object represented by SVG codes, while the other four question types focus on detailed, local understanding of the object. “Color” is color-related questions about specific object parts, which evaluates the localization of the corresponding semantic part. “Count” is about counting the occurrences of certain patterns or semantic parts. “Shape” is about the shape of certain parts of the object, either finding the exact definition of the geometric shape or finding geometric shapes that resemble the object part. Figure 9 gives some SVG examples.
@@
Experimental results and discussion. We find that graphics program understanding, as we operationalize it here, is challenging. The average accuracy of all models (proprietary and open-source) is below 70% (ranging from 30% to 67%) on SVG and below 75% (ranging from 28% to 74%) on CAD. Among these, SVG makes it more difficult for the models to understand as these 2D graphics contain richer semantics. Significant performance improvements are observed in line with scaling laws [113], as larger model sizes consistently lead to gains across various open-source LLMs. For example, Llama-3’s score is improved from 0.429 to 0.548 on SVG and from 0.633 to 0.694 when its size increased from 8B to 70B, Qwen-1.5’s from 0.376 to 0.499 on SVG and 0.486 to 0.632 on CAD with the size from 7B to 110B. We also notice consistent improvements given the same model size and model family but from different generations. For example, Qwen-72B from 0.466 to 0.537, Llama-8B from 0.429 to 0.465 and Llama-70B from 0.548 to 0.574. The consistent performance gain on both the SVG and CAD indicates that the semantic understanding of symbolic graphics programs, including both 2D and 3D understanding, is a fundamental capability that is aligned with the scaling law of LLMs, even if these models have not been specifically trained for this task.
@@ 4.3 Benchmarking Semantic Consistency
Figure 10: The semantic consistency test assesses if semantic understanding remains the same when the program is perturbed without semantically changing its rendered content. Image perturbations result in significant code-level changes because symbolic programs rely on absolute coordinates.
Large language models (LLMs) are exposed to vast amounts of online data, including open-access SVG data. To investigate whether their semantic understanding ability is due to potential data leakage, we propose a semantic consistency test by introducing global translations or rotations to SVG graphics, ensuring SE(2) invariance. Such spatial interventions greatly alter the code representation, as SVG graphics consist of lines and Bezier curves with anchor points, and SE(2) operations change all numerical values in the code. However, the SVG’s semantics—such as shape or color—remain unaffected by this perturbation. This allows us to examine how LLMs behave when the same vector graphics are presented with drastic code-numerical changes (see Appendix A.1 for details). Note that we perform non-trivial coordinate-level perturbations to the code, rather than using SVG transformation functions, to prevent shortcut learning by LLMs. Due to the nested structure of the tested SVG code, we visually inspect the perturbed renderings to ensure that the semantics remain unchanged after perturbation. If the model maintains the performance consistency under these perturbations, it suggests that its semantic understanding is based on a fundamental level of comprehension or visual imagery rather than a trivial memorization.
@@
Evaluation. We measure the semantic consistency with two metrics: 1) the average accuracy of all perturbed SVG inputs’ question-answering accuracy, showing the overall accuracy once the samples are intervened; and 2) the proposed “consistency score” that first counts the most common selection for each group of perturbed samples (where they were translated or rotated from the same SVG program), then divided by the number of perturbations. This score indicates how much the LLMs being consistent (no matter right or wrong) regardless of the drastic program change. If the score is close to 1, it means all the predictions are the same even with totally different input codes, vice versa.
Experimental results and discussion. Our experiments with the SVG-Invariance benchmark demonstrate that most LLMs exhibit robust semantic understanding of graphics programs under translation (T) and translation + rotation (SE(2)) perturbations. In Table 1, most of the LLMs achieve over 80% consistency in both perturbation settings, with half of the models exceeding 90% consistency. Not only do the models remain consistent in their predictions under perturbations, but their performance on perturbed inputs also shows minimal fluctuation compared to their performance on the SVG-Understanding benchmark. We posit that this indicates that the semantic understanding ability that we evaluate of LLMs is unlikely due to data leakage, but rather, could stem from a potential foundational capability to interpret the semantics of deterministic, symbolic graphics programs.
@@ 6 SIT Can Improve General Reasoning Ability
Since we show in Figure 4 that the ability to understand symbolic graphics programs is associated with some fundamental reasoning abilities of LLMs, we are particularly interested in whether symbolic graphcis programs can be used as a novel data source for building better instruction tuning datasets, which can help to improve the general reasoning ability of LLMs. To verify this, we test the instruction-tuned models on a variety of popular LLM benchmarks, including benchmarks focusing on natural language understanding (XNLI [17], IFEval [119], HellaSwag [117], C-Eval [39], CoQA [80], MMLU [34], SQuAD2.0 [77]), generic reasoning (BigBenchHard [92], PIQA [7], AGIEval [118]) and mathematical reasoning (Arithmetic [9], MathQA [3], GSM8k [14], ASDiv [71]).
@@ 7 A critical View on current LLM’s Capability
8 Related Work and Concluding Remarks
@@
Concluding Remarks. We propose a new task over which to evaluate LLMs to contribute towards a multi-faceted assessment of their capabilities: the task of “understand” images – without image; that is, directly from symbolic graphics programs. For such a task, we create SGP-Bench, a benchmark that shows distinguishable results between LLMs and introduce the method of Symbolic Instruction Finetuning (SIT) that improves LLMs’ capability of understanding graphics programs. More importantly, SIT also helps to improve general reasoning ability of LLMs.

### v4

@@ 1 Introduction
  • To improve the symbolic program understanding, we collect an instruction-following dataset and propose symbolic instruction tuning, which can also improve general reasoning performance.
  • Finally, we introduce a symbolic MNIST dataset where the symbolic program is so challenging for LLMs to understand that GPT-4o can only achieve a chance-level performance, while the rendered image is easily recognizable by humans.
@@ 3 Why is Understanding Symbolic
To showcase why semantic understanding of symbolic graphics problems require multiple sophisticated reasoning abilities, we provide a few qualitative examples of LLM’s output in Figure 3 and Figure 4. Figure 3 shows a qualitative example of how OpenAI-o1 reasons over the CAD program. The reasoning process is highly nontrivial, as it requires multiple reasoning abilities, such as numeric perception, spatial reasoning, geometric understanding, long-range planning and common sense. This is also part of the reason that SGP-Bench can well evaluate the general reasoning ability of LLMs. In Figure 4, we query different LLMs (from weak to strong: Llama-3.1-8B, Llama-3.1-70B, OpenAI-o1) by asking which digit this given SVG program represents. We can observe that all LLMs start to reason from the low-level curves and gradually build up its understanding from local components to semantic elements. Specifically, LLMs understand the symbolic program through line-by-line reasoning and then combine the results into an overall semantic understanding. This process is intriguing because it shows that LLMs understand symbolic programs through reasoning rather than memorization. More interestingly, the more powerful LLMs, e.g., OpenAI-o1, show better general understanding of the symbolic program and its fine-grained grounding, which is consistent with the results on SGP-Bench.
@@
To highlight the importance of symbolic graphics programs, we give another example in Figure 5, where a powerful LLM like OpenAI-o1 can suffer from the visual spurious correlation, and in contrast, this spurious correlation can be avoided when using the symbolic graphics program as the visual representation. Specifically, we construct a visual example that resembles the Ebbinghaus illusion, but the conclusion differs from the classic Ebbinghaus illusion (i.e., two orange circles are of the same size, although they look different) as we intentionally make one of the orange circle obviously larger than the other (shown in Figure 5). Then we feed this curated example to OpenAI-o1 and ask which orange circle is larger. We compare three cases: (a) image input; (b) symbolic program input; and (c) image input with indirect symbolic program prompting. Once we first ask the LLM to translate the image into symbolic program, the LLM will no longer suffer from the spurious correlation. The same phenomenon also appears for the Müller-Lyer illusion using Claude-3.5-Sonnet (Figure 6). These examples validate the advantages of reasoning the visual world through symbolic graphics programs.
@@ 4.2 Benchmarking Semantic Understanding
SVG dataset statistics. We collect $1,085$ SVG programs covering $19$ categories, and each program has $4$ semantic multiple-choice questions (with $4$ options), resulting in a total of $4,340$ questions. We ensure that answers are evenly distributed across options. Dataset statistics are given in Figure 8(a). Our SVG benchmark consists of $5$ types of questions, including “Semantic”: $1,085$ questions, “Color”: $864$ questions, “Shape”: $1,217$ questions, “Count”: $819$ questions, and “Reasoning”: $355$ questions. “Semantic” tests the global semantic meaning of the object represented by SVG codes, while the other four question types focus on detailed, local understanding of the object. “Color” is color-related questions about specific object parts, which evaluates the localization of the corresponding semantic part. “Count” is about counting the occurrences of certain patterns or semantic parts. “Shape” is about the shape of certain parts of the object, which is to find geometric shapes that closely resemble the object part. Figure 9 gives some SVG examples.
@@
Experimental results and discussion. We find that graphics program understanding, as we operationalize it here, is challenging. The average accuracy of all models (proprietary and open-source) is below 70% (ranging from 30% to 67%) on SVG and below 75% (ranging from 28% to 74%) on CAD. Among these, SVG makes it more difficult for the models to understand as these 2D graphics contain richer semantics. Significant performance improvements are observed in line with scaling laws [113], as larger model sizes consistently lead to gains across various open-source LLMs. For example, Llama-3’s score is improved from 0.429 to 0.548 on SVG and from 0.633 to 0.694 when its size increased from 8B to 70B, Qwen-1.5’s from 0.376 to 0.499 on SVG and 0.486 to 0.632 on CAD with the size from 7B to 110B. We also notice consistent improvements given the same model size and model family but from different generations. For example, Qwen-72B from 0.466 to 0.537, Llama-8B from 0.429 to 0.465 and Llama-70B from 0.548 to 0.574. The consistent performance gain on both SVG and CAD indicates that semantic understanding of symbolic graphics programs is a fundamental capability that is aligned with the scaling law of LLMs.
@@ 4.3 Benchmarking Semantic Consistency
Figure 10: The semantic consistency test assesses if semantic understanding remains the same when the program is perturbed without semantically changing its rendered content. Image perturbations result in significant code-level changes, as symbolic programs use absolute coordinates.
LLMs are exposed to vast amounts of online SVG data. To investigate whether their semantic understanding ability is due to potential data leakage, we propose a semantic consistency test by introducing global translations or rotations to SVG graphics, ensuring SE(2) invariance. Such spatial interventions greatly alter the code representation, as SVG graphics consist of lines and Bezier curves with anchor points, and SE(2) operations change all numerical values in the code. However, the SVG’s semantics—such as shape or color—remain unaffected by this perturbation. This allows us to examine how LLMs behave when the same vector graphics are presented with drastic code-numerical changes (see Appendix A.1). We perform non-trivial coordinate-level perturbations to the code, rather than using SVG transformation functions, to prevent shortcut learning by LLMs. Due to the nested structure of the tested SVG code, we visually inspect the perturbed renderings to ensure that the semantics remain unchanged after perturbation. If the model performs consistently under these perturbations, it suggests that its semantic understanding stems from a fundamental level of comprehension rather than trivial memorization.
@@
Evaluation. We measure the semantic consistency with two metrics: 1) the average accuracy of all perturbed SVG inputs’ question-answering accuracy, showing the overall accuracy once the samples are intervened; and 2) the proposed “consistency score” that counts the average frequency of the most selected answer to each question for all groups of perturbed samples (where they were translated or rotated from the same SVG program). This score indicates how much the LLMs being consistent (no matter right or wrong) regardless of the drastic program change. If the score is close to 1, it means all the predictions are the same even with totally different input codes.
TED T Cons. SE(2) Cons.
5-10 0.890 0.833
20-25 0.882 0.822
Table 2: Consistency with varying tree edit distance (TED) between original and modified codes.
Experimental results and discussion. Our experiments with the SVG-Invariance benchmark demonstrate that most LLMs exhibit robust semantic understanding of graphics programs under translation (T) and translation + rotation (SE(2)) perturbations. In Table 1, most of the LLMs achieve over 80% consistency in both perturbation settings, with half of the models exceeding 90% consistency. Not only do the models remain consistent in their predictions under perturbations, but their performance on perturbed inputs also shows minimal fluctuation compared to their performance on the SVG-Understanding benchmark. We posit that this indicates that the semantic understanding ability that we evaluate of LLMs is unlikely due to data leakage, but rather, could stem from a potential foundational capability to interpret the semantics of deterministic, symbolic graphics programs. Additionally, we assess the structural alterations introduced by our perturbation operation by calculating the tree edit distance between the original and perturbed code. Our findings indicate that the perturbation leads to varying levels of structural changes in the code. However, we observe no significant correlation between the degree of structural modification and the consistency performance (see Table 2).
@@ 6 SIT Can Improve General Reasoning Ability
Since Figure 4 shows that the ability to understand symbolic graphics programs is associated with some fundamental reasoning abilities of LLMs, we are interested in whether symbolic graphics programs can be used as a novel data source for building better instruction tuning datasets, which can help to improve the general reasoning ability of LLMs. To verify this, we test the instruction-tuned models on a variety of popular LLM benchmarks, including benchmarks focusing on natural language understanding (XNLI [17], IFEval [119], HellaSwag [117], C-Eval [39], CoQA [80], MMLU [34], SQuAD2.0 [77]), generic reasoning (BigBenchHard [92], PIQA [7], AGIEval [118]) and mathematical reasoning (Arithmetic [9], MathQA [3], GSM8k [14], ASDiv [71]).
@@ 7 A critical View on current LLM’s Capability
8 Related Work and Acknowledgment
@@
Acknowledgement. The authors would like to thank Yao Feng and Yandong Wen for helpful suggestions. Additionally, HF would like to thank Hanqi Zhou for her support throughout the project, during which time they became engaged (Fun fact: HF and Hanqi Zhou got married a few hours before the submission deadline). The diamond ring featured in Figure 1 symbolizes this joyous personal milestone and is courtesy of the entire team. This work was supported in part by the German Federal Ministry of Education and Research (BMBF): Tubingen AI Center, FKZ: 01IS18039B, and by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645. WL was supported by the German Research Foundation (DFG): SFB 1233, Robust Vision: Inference Principles and Neural Mechanisms, TP XX, project number: 276693517. KMC acknowledges support from the Marshall Scholarship and Cambridge Trust. AW acknowledges support from a Turing AI Fellowship under grant EP/V025279/1, The Alan Turing Institute, and the Leverhulme Trust via CFI. MJB has received research gift funds from Adobe, Intel, Nvidia, Meta/Facebook, and Amazon. MJB has financial interests in Amazon, Datagen Technologies, and Meshcapade GmbH. While MJB is a consultant for Meshcapade, his research in this project was performed solely at, and funded solely by, the Max Planck Society.